OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ = five sixths. What is the value of cot θ? I need help, I don't understand the steps

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{5}{6}\) all that is missing is the third side, which you find via pythagoras once you have all three sides, you can find any ratio that you like

OpenStudy (anonymous):

cotangent will be "adjacent over opposite"

OpenStudy (anonymous):

so i just times the 5 again and o the same to the 6 to get what it equals?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

would it be 25 & 36?

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

\[a^2+5^2=6^2\] by pythagoras so \[a^2=6^2-5^2=36-25=11\] making \(a=\sqrt{11}\)

OpenStudy (anonymous):

I meant to say that but it came out wrong, so now would I have to do \[\sqrt{11}/5\]